For the reaction 2NO(g) + 2H2(g) yields N2(g) +2H2O(g), the equilibrium constant is 16.7/M. What is the concentration of H2O if NO = 0.34M, H2 = 0.23 M and N2 = 0.40M.


1
Expert's answer
2021-06-02T06:30:56-0400

"K_{eq}=\frac{[N_2][H_2O]^2}{[NO]^2[H_2]^2} \\

16.7 = \frac{0.40 \times [H_2O]^2}{0.34^2 \times 0.23^2} \\

16.7 = \frac{0.40 \times [H_2O]^2}{0.00611524} \\

16.7 = 65.41[H_2O]^2 \\

[H_2O]^2 = \frac{16.7}{65.41} =0.2553 \\

[H_2O] = 0.50"

Answer: 0.50 M